OpenStudy (braydonlevi99):

Solve:

OpenStudy (braydonlevi99):

\[|x-4| + 9 = 3\]

pooja195 (pooja195):

Any ideas on where to start?

OpenStudy (braydonlevi99):

Doesn't it turn to \[|x+4| +9 = 3\]

pooja195 (pooja195):

Nope,try again. We want to isolate the x so we would subtract 9 from both sides to begin with . \[\huge~\rm~\bf~|x+4|+9-9=3-9\] what would we be left with?

OpenStudy (braydonlevi99):

\[|x+4| = -6\] ?

OpenStudy (braydonlevi99):

Then add the 4?

OpenStudy (braydonlevi99):

x = -2 ?

pooja195 (pooja195):

No. It would be |x−4|=-6 and we can't do anything else with it because there's a negative solution. Absolute values cannot have a negative solution (a value less than 0)

OpenStudy (braydonlevi99):

I'm so freaking confused.

pooja195 (pooja195):

On what part?

OpenStudy (braydonlevi99):

All

OpenStudy (braydonlevi99):

You've confused me.

OpenStudy (phi):

do you know what | -9 | simplifies to ?

OpenStudy (braydonlevi99):

To 9

OpenStudy (phi):

what about | -1 | ?

OpenStudy (braydonlevi99):

1

OpenStudy (phi):

now think about this: if you put a negative number inside the | | you get a positive number when you simplify. sound right ?

OpenStudy (phi):

for example | -1 | =1 |-2| = 2 and so on...

OpenStudy (phi):

and if you put in a positive number example: | 9| = 9 we get a positive number after simplifying. in other words, no matter what we put inside the | |, we get out a positive number and that means *it is not possible* for | any number you can think of | = -6 it can't happen.

OpenStudy (phi):

pooja was explaining there is no solution to your equation.